OpenStudy (anonymous):

Which set of side lengths will not form a right triangle? 5, 12, 13 7, 24, 25 9, 16, 20 8, 15, 17

OpenStudy (cwrw238):

take the longest side, square it and see if this equals sum of squares of the other 2 sides (pythagoras) first one forms a right angles triangle because 13^2 # 169 and 5^2 + 12^2 = 25 + 144 = 169

OpenStudy (anonymous):

9, 16, 20 The others are all sets that do. You will know them by heart by the end of Geometry.

OpenStudy (anonymous):

25+144= 169 49+576 = 625 64+ 225 = 289 (That is why the others work)